OpenStudy (anonymous):

Please help me.... Solve for x (2x3 + 4x2 - 32x + 18) ÷ (x - 3)

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

It's not an equation—"solve for x" makes no sense. Are you sure you have copied everything correctly?

OpenStudy (anonymous):

Simplifying 2x4 + 2x3 + -36x2 + -32x + 64 = 0 Reorder the terms: 64 + -32x + -36x2 + 2x3 + 2x4 = 0 Solving 64 + -32x + -36x2 + 2x3 + 2x4 = 0 Solving for variable 'x'. Factor out the Greatest Common Factor (GCF), '2'. 2(32 + -16x + -18x2 + x3 + x4) = 0 Ignore the factor 2. Subproblem 1 Set the factor '(32 + -16x + -18x2 + x3 + x4)' equal to zero and attempt to solve: Simplifying 32 + -16x + -18x2 + x3 + x4 = 0 Solving 32 + -16x + -18x2 + x3 + x4 = 0 The solution to this equation could not be determined. This subproblem is being ignored because a solution could not be determined. The solution to this equation could not be determined.